OpenStudy (anonymous):

MULTI STEP PROBLEM! IT'S A BIGGY!!! Sam solved the equation 3x + 2 = -5(x – 2) incorrectly. Examine Sam’s work and describe his mistake in words. Step 1: 3x + 2 = -5x + 10 Step 2: -2x = 8 Step 3: x = -4 What is the correct solution to 3x + 2 = -5(x – 2) ?

OpenStudy (anonymous):

never mind... this is super easy i don't need any help

OpenStudy (anonymous):

if that is the case...you might as well close this problem

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

by the way....what did you get...I am curious

OpenStudy (anonymous):

Well, I have to solve it first now don't I. I'm doing that right now.

OpenStudy (anonymous):

It is x=1

OpenStudy (anonymous):

3x + 2 = - 5(x - 2) 3x + 2 = - 5x + 10 3x + 5x = 10-2 8x = 8 x = 1 you got it...way to go :)